# Foregrounding Artist Opinions: A Survey Study on Transparency, Ownership, and Fairness in AI Generative Art

### v4

@@ 3.1. Survey Methodology
Our anonymous original survey of artists is an observational study (N=516) of artists over the age of 18 years that was collected using a convenience sample in phase 1 and supplemented with a Qualtrics survey panel in phase 2 (demographic breakdown can be seen in Table 1. Qualtrics is a reputable academic survey panel platform that has been shown to yield reliable results in comparison to other survey platforms (Boas et al., 2018). For comparison of the distributions of the US art workforce) based on the demographics is presented in Table 4. Incomplete surveys were dropped from the analysis, leaving a final N of 459. Participants provided consent by agreeing to the survey information sheet. They were then prompted to respond to inquiries regarding their artistic identity and financial interactions within the art domain. Participants were presented with a brief introduction to AI-generated art, including an illustrative example, and were asked to indicate their familiarity with AI art tools and models. Subsequently, participants were queried on their perceptions regarding ownership of art produced by AI systems using a 5-point Likert scale. Further questions focused on artists’ preferred compensation criteria for their work being used in AI-generated art models. Participants were also asked to evaluate the impact of AI art models on their field, their perception of these models as potential threats, and whether they believed creators of such models should disclose their data sources using a 5-point Likert scale. Finally, demographic information about the participants’ backgrounds was collected. An appendix that enumerates all survey questions is available in Appendix I. This survey was approved as “exempt” by the University of Vermont Institutional Review Board (IRB).
@@
Phase 2 Sampling: Phase 2 (n = 264) was collected via a targeted survey panel conducted by Qualtrics from September 12-23, 2023 on the Qualtrics online survey platform.
3.2. Data
Our original survey of artists is a convenience survey (N=459).We transformed all survey response variables of interest into numerical form to analyze our survey results (survey questions 8,9,10,13,14,15 as seen in Appendix I). All Likert survey questions were converted from ‘Strongly Agree,’ ‘Agree,’ ‘Neutral,’ ‘Disagree,’ and ‘Strongly Disagree’ to an ordinal scale of 1,2,3,4,5. For our ordinal logistic regression analysis, we combine ‘Strongly Agree’ and ‘Agree’ to be equal to 3, ’Neutral’ to be equal to 2, and ‘Disagree’ and ‘Strongly Disagree’ to be equal to 1. In the original logistic regression, we reduced the dimensionality of the Likert scale to investigate the positive, neutral, and negative sentiment of the response questions in relation to our secondary variables.
All nominal and categorical variables (survey questions 1,2,3,4,5,11,12,16,17,18,19, as seen in Appendix I) were transformed into binary variables. Survey questions that included an “other” option or write-in answers were also categorized (e.g., in question 2, we added an additional art practice, which included digital art from the write-in answers).
Demographic questions (survey questions 16, 17, 18, 19 as seen in Appendix I) included asking survey participants their gender identity to identify whether they were white or a person of color, choosing their age range and country of residence. All demographic questions were transformed into binary variables.
3.3. Analytical Methods
@@
where p is the probability of success. Unlike probability, odds can be from 1 to positive infinity (Bruin, 2011a). The odds ratio (OR, or ) is simply the odds of one event divided by the odds of another. The OR can be from 0 to positive infinity, with an OR equal to 1 indicating equal odds of success between events. The natural logarithm of the OR is often used, as the transformation renders the distribution symmetric about 0 (Agresti, 2012).
Ordinal Logistic Regression Ordinal logistic regression (OLR) is a generalization of regression (either multiple linear or binomial logistic) used for an ordered categorical response variable and any number of categorical or numeric secondary variables. OLR finds the logit (also known as log-odds) of the dependent variable (Y) being less than or equal to a certain level (denoted j) (Bruin, 2011b). which can be written as:
@@ 4. Results
This research question investigates artists’ perceptions of Generative AI models as potential threats to art workers. Our findings reveal that 61.87% of survey participants hold the view that Generative AI models pose a threat to art workers. Thus, while a majority of artists perceive AI as a threat, it is not an overwhelming majority. This finding is related to survey question no. 15 (see Appendix I for survey questions).
An ordinal logistic regression (OLR) analysis was performed to examine the relationship between the Likert scale responses to RQ1 (dependent variable) and the secondary variables (all odds ratios can be seen in Appendix III), focusing on the odds ratios derived from the OLR model (as seen in Figure 2).
The results demonstrate a significant relationship (p-value 0.05, all detailed odds ratios can be seen in Appendix III) between agreeing (represents a combination of both the responses ‘Agree’ and ‘Strongly Agree’ in the Likert scale) that Generative AI models are a threat to art workers and a subset of our secondary variables. Here, we see that artists ages 18-29 are more likely to agree that there is a threat and that artists who have used Generative AI art models to create art are less likely to think that it is a threat to the art workforce.
@@
This research question investigates artists’ perceptions of AI models as positive development to the field of art. We ran the same analysis for our second research question and found that 44.88% of the survey participants agree that Generative AI art models are a positive development in the field of art. Notably, 27.02% of artists selected ’neutral’ on the Likert scale for this question. This analysis pertains to survey question number 14 (see Appendix I for survey questions).
We conducted an ordinal logistic regression (OLR) analysis to investigate the association between Likert scale responses to RQ2 (dependent variable) and the secondary variables with a focus on the odds ratios obtained from the OLR model (as seen in Figure 3). Our analysis indicates a strong relationship (p-value 0.05, all OLR detailed results can be seen in Appendix III) between agreeing that Generative AI art models are a positive development in the field of art and a subset of our secondary variables. Here, we see that 30-49 years olds, males, persons of color, participants who have used AI art models, and participants who said they would donate their artwork to train models were more likely to agree that generative AI models are a positive development for the field of art. Whereas makers (note that makers were a small n (n=5) with a fairly large range in the confidence interval bars represented in Figure 3), females, and those who identify as artists were less likely to agree that AI models are a positive development for the field of art.
@@
This research question investigates artists’ perceptions of what types of disclosure are fair when using their artwork to train an AI art model. Our results show a consensus among a large majority of artists that transparency and disclosure regarding the specific images used in training an AI model are essential prerequisites for them to deem the model fair and acceptable. We see from our survey results that 80.17% of the survey participants agree that creators of AI models should disclose the art and images used to train their models in detail. The analysis for RQ3 is related to survey question 13 (all survey questions can be seen in Appendix I), which we used to run an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ3 and our secondary variables. The results of the OLR, displayed in Figure 4 as the odds ratio, show a strong relationship (p-value 0.05, all odds ratios can be seen in Appendix III) between agreeing with the disclosure requirement and a subset of our secondary variables.
@@
Finally, we look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. The analysis for RQ4 is related to survey questions No. 8, 9, and 10 (all survey questions can be seen in Appendix I).
@@
To investigate the relationship between the Likert scale responses to RQ4B and the secondary variables, we conducted an ordinal logistic regression (OLR), presented in Figure 6 displaying the odds ratios derived from the analysis. Our results show that there is a strong relationship (p-value 0.05, all detailed odds ratios can be seen in Appendix III) between agreeing that the artist whose style is recognizable in the output should be the owner of the art and a subset of our secondary variables. We see that participants who are persons of color are more likely to agree that the owner of the Generative AI artwork should be the artist whose style is recognizable and mimicked by the Generative AI art model. Whereas participants who have used Generative AI to make art were less likely to agree that the owner of the Generative AI artwork should be the artist whose style is recognizable and mimicked by the Generative AI art model.
@@
Our analysis suggests a strong relationship (p-value 0.05, all odds ratios can be seen in Appendix III) between agreement on AI companies owning the resulting artwork and a subset of our secondary variables. Participants who considered themselves tattooists (note that tattooists were a small n (n=9) with a fairly large range in the confidence interval bars represented in Figure 7), males, persons of color, or those who said they would donate their artwork to train generative AI models were more likely to agree that the owner of the Generative AI artwork should be the AI model creators. Whereas participants with a general familiarity with Generative AI Art, females, and non-binary participants were less likely to agree that the owner of the Generative AI artwork should be the AI model creators.
5. Discussion
The impact of Generative AI on artists is a complex global issue with impacts in many arenas, and it must be investigated through further multidisciplinary collaboration. We hope this work contributes to the existing literature on the ethics of AI-generated art and helps to provide a voice for artists on this topic.
@@ 5. Discussion
6. Acknowledgements
We are thankful to the anonymous reviewers for their insightful comments. Institutional Review Board Approval: The survey in this project is CHRBSS (Behavioral) STUDY00002379, approved by the University of Vermont I.R.B. on 3/14/2023. The authors would like to thank Maria Sckolnick, John Meluso, Kendall Fortney, Randall Harp, Laurent Hébert-Dufresne, Jean-Gabiel Young, and Winn Gillen, for their comments on early versions of this work. The Alfred P. Sloan Foundation, The National Science Foundation award #2242829, and MassMutual under the MassMutual Center of Excellence in Complex Systems and Data Science support this work. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the aforementioned financial supporters.
Appendix A Appendix I: Survey Questions
@@
Appendix B Appendix II: Full table of descriptive statistics
@@
Appendix C Appendix III: All tables for ordinal logistic regressions
@@
Appendix D Appendix IV: Contingency Tables
Appendix E Appendix V: Tests for Independence
E.0.1. Kruskal-Wallis test for independence:
@@
Kruskal–Wallis Test The Kruskal-Wallis H-test is a non-parametric test that extends the Mann-Whitney U-test to compare differences between groups for an ordinal or continuous dependent variable. The explanatory groups may be ranked or unranked (website, 2018). Its null hypothesis is stochastic homogeneity, which is equivalent to equality of the expected value of rank means, conditional on the homogeneity of rank variances (Vargha and Delaney, 1998).
E.0.2. Cochran-Armitage test for independence:
@@
Cochran–Armitage Test The Cochran-Armitage test for trend uses weighted linear regression coefficients to test for trends in proportion between ordered levels of the response variable for a binary response variable. Its null hypothesis is a lack of trend, meaning that the proportions of each outcome are the same for each explanatory level (Iskander et al., 1983).

### v5

@@ Methods
Our anonymous original survey of artists is an observational study (initial N=516 (includes incomplete answers), final N=459 (only complete answers)) of artists over the age of 18 years that was collected using a convenience sample in phase 1 and supplemented with a Qualtrics survey panel in phase 2 (demographic breakdown can be seen in Table 1. Qualtrics is a reputable academic survey panel platform that has been shown to yield reliable results in comparison to other survey platforms (Boas, Christenson, and Glick 2018). A comparison of the distributions of the US art workforce) based on the demographics is presented in Appendix B. Survey participants were recruited based on their self-identification as artists and must be over the age of 18. All survey questions are original and developed by the authors. Incomplete surveys were dropped from the analysis, leaving a final N of 459. Participants provided consent by agreeing to the survey information sheet. They were then prompted to respond to inquiries regarding their artistic identity and financial interactions within the art domain. Participants were presented with a brief introduction to AI-generated art, including an illustrative example, and were asked to indicate their familiarity with AI art tools and models. Subsequently, participants were queried on their perceptions regarding ownership of art produced by AI systems using a 5-point Likert scale. Further questions focused on artists’ preferred compensation criteria for their work being used in AI-generated art models. Participants were also asked to evaluate the impact of AI art models on their field, their perception of these models as potential threats, and whether they believed creators of such models should disclose their data sources using a 5-point Likert scale. Finally, demographic information about the participants’ backgrounds was collected. An appendix that enumerates all survey questions is available in Appendix A, which can be found in our supplementary materials (Lovato et al. 2024). This survey was approved as “exempt” by the University of Vermont Institutional Review Board (IRB).
Sampling limitations: In future work, we would like to sample a much bigger representative sample that specifically targets artistic practices and demographics. In addition, we are interested in how the dynamics of these opinions change over time with surveys that assess artist opinions at regular intervals.
@@
Phase 2 Sampling: Phase 2 (n = 264) was collected via a targeted survey panel conducted by Qualtrics from September 12-23, 2023 on the Qualtrics online survey platform. The inclusion criteria for the study in Phase 2 were the same as in Phase 1.
Data
Our original survey of artists is an observational survey that took place over two sample phases (total N=459 completed surveys over both phases). We transformed all survey response variables of interest into numerical form to analyze our survey results (survey questions 8,9,10,13,14,15 as seen in Appendix A). All Likert survey questions were converted from ‘Strongly Agree,’ ‘Agree,’ ‘Neutral,’ ‘Disagree,’ and ‘Strongly Disagree’ to an ordinal scale of 1,2,3,4,5. For our ordinal logistic regression analysis, we combine ‘Strongly Agree’ and ‘Agree’ to be equal to 3, ’Neutral’ to be equal to 2, and ‘Disagree’ and ‘Strongly Disagree’ to be equal to 1. In the original logistic regression, we reduced the dimensionality of the Likert scale to investigate the positive, neutral, and negative sentiment of the response questions in relation to our secondary variables.
All nominal and categorical variables (survey questions 1,2,3,4,5,11,12,16,17,18,19, as seen in Appendix A) were transformed into binary variables. Survey questions that included an “other” option or write-in answers were also categorized (e.g., in question 2, we added an additional art practice, which included digital art from the write-in answers).
Demographic questions (survey questions 16, 17, 18, 19 as seen in Appendix A) included asking survey participants their gender identity to identify whether they were white or a person of color, choosing their age range and country of residence. All demographic questions were transformed into binary variables.
Analytical Methods
@@
where p is the probability of success. Unlike probability, odds can be from 1 to positive infinity (Bruin 2011a). The odds ratio (OR, or ) is simply the odds of one event divided by the odds of another. The OR can be from 0 to positive infinity, with an OR equal to 1 indicating equal odds of success between events. The natural logarithm of the OR is often used, as the transformation renders the distribution symmetric about 0 (Agresti 2012).
Ordinal Logistic Regression Ordinal logistic regression (OLR) is a generalization of regression (either multiple linear or binomial logistic) used for an ordered categorical response variable and any number of categorical or numeric secondary variables. OLR finds the logit (also known as log-odds) of the dependent variable (Y) being less than or equal to a certain level (denoted j) (Bruin 2011b). which can be written as:
@@ Results
This research question investigates artists’ perceptions of Generative AI models as potential threats to art workers. Our findings reveal that 61.87% of survey participants hold the view that Generative AI models pose a threat to art workers. Thus, while a majority of artists perceive AI as a threat, it is not an overwhelming majority. This finding is related to survey question no. 15 (see Appendix A for survey questions).
An ordinal logistic regression (OLR) analysis was performed to examine the relationship between the Likert scale responses to RQ1 (dependent variable) and the secondary variables (all odds ratios can be seen in Appendix C), focusing on the odds ratios derived from the OLR model (as seen in Figure 2).
The results demonstrate a significant relationship (p-value 0.05, all detailed odds ratios can be seen in Appendix C) between agreeing (represents a combination of both the responses ‘Agree’ and ‘Strongly Agree’ in the Likert scale) that Generative AI models are a threat to art workers and a subset of our secondary variables. Here, we see that artists ages 18-29 are more likely to agree that there is a threat and that artists who have used Generative AI art models to create art are less likely to think that it is a threat to the art workforce.
@@
This research question investigates artists’ perceptions of AI models as positive development to the field of art. We ran the same analysis for our second research question and found that 44.88% of the survey participants agree that Generative AI art models are a positive development in the field of art. Notably, 27.02% of artists selected ’neutral’ on the Likert scale for this question. This analysis pertains to survey question number 14 (see Appendix A for survey questions).
We conducted an ordinal logistic regression (OLR) analysis to investigate the association between Likert scale responses to RQ2 (dependent variable) and the secondary variables with a focus on the odds ratios obtained from the OLR model (as seen in Figure 3). Our analysis indicates a strong relationship (p-value 0.05, all OLR detailed results can be seen in Appendix C) between agreeing that Generative AI art models are a positive development in the field of art and a subset of our secondary variables. Here, we see that 30-49 years olds, males, persons of color, participants who have used AI art models, and participants who said they would donate their artwork to train models were more likely to agree that generative AI models are a positive development for the field of art. Whereas makers (note that makers were a small n (n=5) with a fairly large range in the confidence interval bars represented in Figure 3), females, and those who identify as artists were less likely to agree that AI models are a positive development for the field of art.
@@
This research question investigates artists’ perceptions of what types of disclosure are fair when using their artwork to train an AI art model. Our results show a consensus among a large majority of artists that transparency and disclosure regarding the specific images used in training an AI model are essential prerequisites for them to deem the model fair and acceptable. We see from our survey results that 80.17% of the survey participants agree that creators of AI models should disclose the art and images used to train their models in detail. The analysis for RQ3 is related to survey question 13 (all survey questions can be seen in Appendix A), which we used to run an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ3 and our secondary variables. The results of the OLR, displayed in Figure 4 as the odds ratio, show a strong relationship (p-value 0.05, all odds ratios can be seen in Appendix C) between agreeing with the disclosure requirement and a subset of our secondary variables.
@@
Finally, we look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. The analysis for RQ4 is related to survey questions No. 8, 9, and 10 (all survey questions can be seen in Appendix A).
@@
To investigate the relationship between the Likert scale responses to RQ4B and the secondary variables, we conducted an ordinal logistic regression (OLR), presented in Figure 6 displaying the odds ratios derived from the analysis. Our results show that there is a strong relationship (p-value 0.05, all detailed odds ratios can be seen in Appendix C) between agreeing that the artist whose style is recognizable in the output should be the owner of the art and a subset of our secondary variables. We see that participants who are persons of color are more likely to agree that the owner of the Generative AI artwork should be the artist whose style is recognizable and mimicked by the Generative AI art model. Whereas participants who have used Generative AI to make art were less likely to agree that the owner of the Generative AI artwork should be the artist whose style is recognizable and mimicked by the Generative AI art model.
@@
Our analysis suggests a strong relationship (p-value 0.05, all odds ratios can be seen in Appendix C) between agreement on AI companies owning the resulting artwork and a subset of our secondary variables. Participants who considered themselves tattooists (note that tattooists were a small n (n=9) with a fairly large range in the confidence interval bars represented in Figure 7), males, persons of color, or those who said they would donate their artwork to train generative AI models were more likely to agree that the owner of the Generative AI artwork should be the AI model creators. Whereas participants with a general familiarity with Generative AI Art, females, and non-binary participants were less likely to agree that the owner of the Generative AI artwork should be the AI model creators.
Discussion
The impact of Generative AI on artists is a complex global issue with consequences in many arenas, and it must be investigated through further multidisciplinary collaboration. We hope this work contributes to the existing literature on the ethics of AI-generated art and helps to provide a voice for artists on this topic.
@@ Discussion
Acknowledgements
We are thankful to the anonymous reviewers for their insightful comments. Institutional Review Board Approval: The survey in this project is CHRBSS (Behavioral) STUDY00002379, approved by the University of Vermont I.R.B. on 3/14/2023. The authors would like to thank Maria Sckolnick, John Meluso, Kendall Fortney, Randall Harp, Laurent Hébert-Dufresne, Jean-Gabiel Young, and Winn Gillen, for their comments on early versions of this work. The Alfred P. Sloan Foundation award #2024-22498, The National Science Foundation award #2242829, and the MassMutual Center of Excellence in Complex Systems and Data Science support this work. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the aforementioned financial supporters.
Appendix A Appendix A: Survey Questions
1.
@@ Appendix A Appendix A: Survey Questions
Appendix B Appendix B: Full Table of Descriptive Statistics
@@
Appendix C Appendix C: All Tables for Ordinal Logistic Regressions
@@
Appendix D Appendix D: Contingency Tables
Appendix E Appendix E: Tests for Independence
Kruskal-Wallis Test for Independence:
@@ Appendix E Appendix E: Tests for Independence
Kruskal–Wallis Test The Kruskal-Wallis H-test is a non-parametric test that extends the Mann-Whitney U-test to compare differences between groups for an ordinal or continuous dependent variable. The explanatory groups may be ranked or unranked (website 2018). Its null hypothesis is stochastic homogeneity, which is equivalent to equality of the expected value of rank means, conditional on the homogeneity of rank variances (Vargha and Delaney 1998).
Cochran-Armitage Test for Independence:
@@
Cochran–Armitage Test The Cochran-Armitage test for trend uses weighted linear regression coefficients to test for trends in proportion between ordered levels of the response variable for a binary response variable. Its null hypothesis is a lack of trend, meaning that the proportions of each outcome are the same for each explanatory level (Iskander, Lakhtakia, and Durney 1983).
@@
Appendix F: Code and Data Availability
All data and all code are available from our GitHub repository, which can be found at the following URL:
https://github.com/juniperlovato/AASP